OpenStudy (anonymous):

At 9 am a car (A) began a journey from a point, traveling at 40 mph. At 10 am another car (B) started traveling from the same point at 60 mph in the same direction as car (A). At what time will car B pass car A? I know this is pretty simple but evry time I look at my answer it looks wrong.

OpenStudy (nathan917):

* After t hours the distances D1 traveled by car A is given by D1 = 40 t * Car B starts at 10 am and will therefore have spent one hour less than car A when it passes it. After (t - 1) hours, distance D2 traveled by car B is given by D2 = 60 (t-1) * When car B passes car A, they are at the same distance from the starting point and therefore D1 = D2 which gives 40 t = 60 (t-1) * Solve the above equation for t to find t = 3 hours * Car B passes car A at 9 + 3 = 12 pm is the answer.

OpenStudy (nathan917):

Does that look wrong? :P

OpenStudy (anonymous):

No, it dosn't my answer kept coming up as 11 pm

OpenStudy (nathan917):

And welcome to open study!

OpenStudy (anonymous):

Thank's